OpenStudy (anonymous):

Essay Question: Show all work. Suzie's flower shop estimates that supplies costs for one month will be given by the polynomial 2x^2-7x-4. She estimates that utilities costs for one month is given by 3x-3. find and simplify an expression that would estimate Suzie's total costs for one month. This is alot like nickie's

OpenStudy (anonymous):

total costs = supplies costs + utilities costs total costs = (x^2 - 3x + 6) + (7x - 9) total costs = x^2 + 4x - 3

OpenStudy (anonymous):

Is it an essay question???? lol

OpenStudy (anonymous):

says that it is

OpenStudy (anonymous):

so x^2+4x-3 is the answer?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

now how did you get this seashell

OpenStudy (anonymous):

total costs = supplies costs + utilities costs total costs = (x^2 - 3x + 6) + (7x - 9) total costs = x^2 + 4x - 3

OpenStudy (anonymous):

this does not look right can you explain how you got what you did with 2x^2-7x-4 and 3x-3 ?

OpenStudy (anonymous):

2x^2-7x-4 +3x-3 =??????? it is the answe mr.seashelllll.